# Network of velocity-coherent filaments formed by supersonic turbulence in a very-high-velocity HI cloud

### v1

@@
The warm neutral medium (WNM) was thought to be subsonically/transonically turbulent, and it lacks a network of intertwined filaments 1, 2 that are commonly seen in both molecular clouds (refs. 3, 4, 5, 6, 7, 8, 9, 10) and cold neutral medium (CNM) (refs.11, 12, 13). Here, we report Hi 21 cm line observations of a very-high-velocity (-330 km s-1 -250 km s-1) cloud (VHVC), using the Five-hundred-meter Aperture Spherical radio Telescope (FAST) with unprecedented resolution and sensitivity. For the first time, such a VHVC is clearly revealed to be a supersonic WNM system consisting of a network of velocity-coherent Hi filaments. The filaments are in the forms of slim curves, hubs, and webs, distributed in different layers within the position-position-velocity (ppv) data cube. The entire cloud has skewed log-normal probability distribution of column density and the filaments themselves show asymmetrical radial density profiles, indicating shock compression by supersonic magnetohydrodynamic (MHD) turbulence, as is also confirmed by our MHD simulation (sonic Mach number and Alfvén Mach number ). This work suggests that hierarchical filaments can be established by shocks in a low-density WNM, where gravity is negligible, offering a viable pathway to structure formation in the earliest evolutionary phases of the interstellar medium (ISM).
@@
Hierarchical density structures in the warm neutral medium (WNM) have not yet been well characterized by observations even though some theoretical works have predicted that elongated low-density parallel filaments can be formed in the sub-Alfvénic WNM 14. High-velocity clouds (HVCs; km s-1) that contain a significant amount of WNM gas and locate in relatively isolated environments provide an excellent laboratory to explore this issue. HVCs are Hi clouds at radial velocities incompatible with the regular rotation of the Galactic disc15, 16, and the absorption lines of distant halo stars indicate that they are generally located several kpc above the Galactic plane (refs.17, 18, 19, 20). The overall negative velocities of HVCs (refs.16, 18, 21) suggest that they may trace the infall of early-phase low-metalicity material onto the plane of the Milky Way (refs.22, 23, 24, 25). HVCs under the view of the all-sky Hi surveys such as HI4PI21 are in the form of atomic gas streams or complexes, because of limited resolution (; see also Extended Data Fig. 1). A few HVCs with km s-1 that were observed by the GALFA-Hi survey of Arecibo 26 have been revealed to be WNM systems consisting of slim filaments27, 28. Most of those filaments are aligned in parallel patterns, and may be regulated by the global dynamics of the gas streams, or by the dynamic processes in the Galactic plane 28, 29. In contrast, the very-high-velocity clouds (VHVCs; km s-1) are for certain not coupled with the Galactic disc, and may have a different origin from common HVCs (refs. 29, 30, 31, 32). Detailed studies of VHVCs, however, are extremely rare in the past. Observations of the Hi 21 cm line (, =1-0; MHz)15 with sufficiently high resolutions to reveal the internal morphology and gas dynamics of VHVCs are essential for studying their origins and the underlying physical processes involved in the formation of hierarchical gas structures within the WNM.
@@
The HI4PI survey (refs.33, 34, 35, 36) provided so far the most detailed Hi images of all-sky HVCs21. VHVCs are dominated by the Magellanic Stream (MS) 37, 38 and the anticentre (AC) complex (refs.39, 40, 29). The VHVC component of the AC complex, denoted as G165 VHVC in this work, is the most notable one in brightness and size on the sky (Extended Data Fig. 1), which stands out as an ideal object for investigating filament formation in extreme environments. Besides the VHVC component (-330 km s-1 -250 km s-1), at the line of sight of G165 there is also an HVC component (-140 km s-1 -80 km s-1) and a foreground component (-20 km s-1 20 km s-1). G165 HVC is part of the Cohen Stream, which was measured to have a distance bracket of 5.0 to 11.7 kpc18. In contrast, the H brightness of G165 VHVC implies a distance () on the order of 15 kpc18, the value we adopted. The G165 VHVC is gravitationally unbound with a size of 100 pc, and a total gas mass of (see methods). Limited by poor angular resolution (), the inner structures of the G165 VHVC were not resolved but did show clumpy features in the HI4PI survey.
@@ A WNM system with supersonic dynamics
We mapped the G165 VHVC using the Five-hundred-meter Aperture Spherical radio Telescope (FAST), with unprecedented angular resolution (), velocity resolution ( km s-1), and line sensitivity (60 mK per channel). G165 VHVC is sufficiently resolved in the integrated intensity map of the Hi 21 cm line (left panel in Fig. 1). Different velocity subcomponents of the G165 VHVC are spectrally blended, but can be well separated by careful Gaussian decomposition (Fig. 1; see Sect. Methods for details). These subcomponents are distributed in several layers in the position-position-velocity (ppv) cube (Fig. 2). The layers are interlaced with each other, and show large-scale velocity fluctuations (or velocity wiggles) as also witnessed in molecular interstellar medium 41. The central velocities () of the subcomponents are divided into five groups ranging from km s-1 to km s-1 (Fig. 3a), with no one-to-one correspondence between the velocity groups and gas layers.
@@
The line widths (full width at half maximum; FWHM) of the velocity subcomponents are slightly larger than 20 km s-1 (Fig. 3c), the value of the typical thermal line widths of the WNM42, 43, 2. Assuming a gas kinetic temperature of the G165 VHVC of 6000 K, the typical value of WNM2, the thermal dispersion is estimated as 7 km s-1, yielding the transonic () nature of individual velocity subcomponents. However, different velocity groups are separated by 20 km s-1, suggesting the whole cloud could be in supersonic turbulence. Considering that different gas layers are intersected, forming a dynamically correlated system (see also the next sections), all velocity groups can be treated as a whole system. The total dispersion of the centroid velocities (), a good tracer of the internal turbulent motions within a cloud, is up to 20 km s-1, indicating a global of 3. We note that the observed distribution of line widths (Fig. 3c) can be well reproduced by our MHD simulation of supersonic and trans-Alfvénic gas ( and ). The same simulation also spans a large velocity range with several velocity groups (Fig. 3b), as the same in observations.
@@
The median values of column density ( cm-2) and volume density (0.2 cm-3) of the G165 VHVC (Sect. Methods) are also similar to the typical values of the WNM42. The column density probability distribution function (-PDF) of the decomposed subcomponents is higher-end suppressed (or lower-end enhanced), and deviates from a lognormal shape (Fig. 3d). The lognormal behavior of an -PDF is usually explained by supersonic turbulence44, 45, 46, but the intermittency of a supersonic system caused by head-on collisions of strongly shocked gas could lead to considerable deviations47, 48. The observed -PDF can also be depicted by our MHD simulation models of supersonic and trans-Alfvénic gas ( and ; see below).
@@
To conclude, we consider the G165 VHVC as a WNM system with supersonic dynamics. This is significantly different from the majority of the WNM in Galactic plane or in HVCs (e.g., Smith cloud), where turbulence is subsonic to transonic 14, 28, 49, 46.
@@
Filamentary structures in the G165 region, which cannot be resolved by HI4PI, are clearly seen by FAST on the Hi moment 0 maps of its VHVC, HVC, and foreground components (Extended Data Fig. 2). The HVC components show parallel filaments elongated along the Cohen Stream, similar to the famous Smith HVC 28. In contrast, both the foreground and VHVC components show web-like filamentary structures. The foreground component has no clear boundary, while the VHVC shows clear boundaries defined by long and slim filaments. Another difference between the foreground and VHVC components is that the filaments tend to be in parallel mode on the channel maps of the foreground component, but stay intersected at different velocity slices of the VHVC components (Extended Data Fig. 2). These comparisons indicate that the filamentary system of the G165 VHVC is really web-like rather than in parallel mode, which is substantially different from those in typical HVCs (e.g., Smith HVC 28) and low/intermediate velocity Hi clouds 11, 13.
@@
Filamentary structures in the G165 VHVC are visible in different layers of the velocity subcomponents (Fig. 2). However, this may be misleading due to projection effects, which enhances the surface intensity at places tangent to the line of sight. We developed a tool, ClimbRidge, based on a density ridge algorithm50, 51, to extract the density ridges of G165 VHVC in both 2-d and 3-d cases (Sect. Methods). ClimbRidge gives similar results as the well known tool FilFinder in 2-d case (Extended Data Fig. 3), and can be naturally applied to 3-d (ppv) data to trace truly velocity-coherent filaments (Fig. 2 and Extended Data Fig. 4). Minimal Spanning Trees (MSTs) of the points on the density ridges were built (Extended Data Fig. 6; see Sect. Methods), and the lengths of filament segments were measured (Fig. 3e). As one can see, the distribution of segment lengths in our MHD simulation of supersonic and trans-Alfvénic gas ( and ) remarkably resembles the observed one.
@@
The filaments in the G165 VHVC are mainly distributed in three layers in ppv space (denoted as L1, L2 and L3 in Extended Data Fig. 4). Each layer has an angular size larger than 2.5∘. On the plane of (RA, DEC), the G165 VHVC can be divided into its eastern, western, and southern parts. L1 is the main body of the western part. L2 spans from the eastern part to the western part, with a global velocity gradient of 0.2 km s-1 per arcmin (0.05 km s-1 pc-1). L3 extends from the western part to the southern part, with a global velocity gradient of 0.3 km s-1 per arcmin (0.07 km s-1 pc-1) (Extended Data Fig. 4). The gas flares of L2, with their velocities spreading to km s-1, also has a filamentary structure. The filaments of the G165 VHVC are intersected with each other, building a filamentary network in the ppv space.
@@
There are long and slim filaments that go through different layers, indicating that different layers are indeed not independent and form a dynamically correlated system. Three representative filaments (denoted as B1, B2, and B3) are highlighted in the right panel of Fig. 2 and Extended Data Fig. 4. We view these long filaments from different viewing angle in the ppv space (Extended Data Fig. 5), confirming that the density ridges do trace the skeletons of the filamentary structures in the ppv space. Basic parameters of the three filaments are fitted and listed in Extended Data Table 1. The velocity gradient of the filaments is comparable with that of the gas layers (0.1 km s-1 pc-1), contributing a velocity span of km s-1. This also suggests that the filaments in G165 VHVC are supersonic. The filaments of the G165 VHVC tend to be radially asymmetrical on the - plane (for example see Extended Data Fig. 5). The intensity difference between the two sides of the filament is comparable with the intensity enhancement of the central ridge (Extended Data Fig. 7). Our MHD simulations suggest that the asymmetrical density enhancement of filaments could be induced by supersonic turbulence (Fig. 3f).
@@
Analog to filaments in Molecular Clouds
Hierarchical filamentary structures are ubiquitous in Galactic molecular clouds (MCs), in the form of slim curves52, hubs53, and webs3. Velocity gradients along the filaments are typically km s-1 pc-1 in MCs54, 7, and can reach 5 km s-1 pc-1 within massive molecular clumps8. It was often interpreted as gas flow along the filament under the effect of gravity. This work implies that velocity gradients along the filaments have been established in early-phase atomic clouds like the G165 VHVC, and will increase from 0.1 km s-1 pc-1 to several km s-1 pc-1 as their natal clouds vary from hundred-pc-scale atomic clouds to sub-pc-scale molecular clumps. The contribution of turbulence to the velocity gradients along the filaments should not be ignored. In spite of the possible different origins of velocity gradients, the filaments of the G165 VHVC could also play a role in transporting material between different layers (Extended Data Fig. 4). The web-like filaments within G165 VHVC do not just intersect with each other. Instead, different filaments can intersect each other constituting hub-like structures with mass assembling at the intersection knots. For example, the long filament B1 and its short fibers form such a hub-filament system (Extended Data Fig. 5). To exclude projection bias, we changed the viewing angles to find out the fibers that are truly connected to B1. Three such subbranches are found, each of which contributes an intensity enhanced hub at the intersection point. In a word, all filamentary types (slim curves, hubs, and webs) that commonly exist in MCs are also revealed in the G165 VHVC, indicating that the origins of the filaments in these two distinctly different types of ISMs may be regulated by some shared physical mechanisms including, most likely, supersonic MHD turbulence. Observations of more VHVCs in the future should yield a deeper understanding of the physical nature of VHVCs and their formation.

### v2

@@
The warm neutral medium (WNM) was thought to be subsonically/transonically turbulent, and it lacks a network of intertwined filaments that are commonly seen in both molecular clouds and cold neutral medium (CNM). Here, we report Hi 21 cm line observations of a very-high-velocity (-330 km s-1 -250 km s-1) cloud (VHVC), using the Five-hundred-meter Aperture Spherical radio Telescope (FAST) with unprecedented resolution and sensitivity. For the first time, such a VHVC is clearly revealed to be a supersonic WNM system consisting of a network of velocity-coherent Hi filaments. The filaments are in the forms of slim curves, hubs, and webs, distributed in different layers within the position-position-velocity (ppv) data cube. The entire cloud has skewed log-normal probability distribution of column density and the filaments themselves show asymmetrical radial density profiles, indicating shock compression by supersonic magnetohydrodynamic (MHD) turbulence, as is also confirmed by our MHD simulation (sonic Mach number and Alfvén Mach number ). This work suggests that hierarchical filaments can be established by shocks in a low-density WNM, where gravity is negligible, offering a viable pathway to structure formation in the earliest evolutionary phases of the interstellar medium (ISM).
@@
Hierarchical density structures in the warm neutral medium (WNM) are not yet well characterized by observations 1, 2, in contrast to the clearer picture emerging for molecular clouds 3, 4, 5, 6, 7, 8, 9, 10 and the cold neutral medium (CNM) 11, 12, 13, even though theory predicts that sub-Alfvénic turbulence can give rise to elongated, low-density filaments in the WNM 14. High-velocity clouds (HVCs; km s-1) that contain a significant amount of WNM gas and locate in relatively isolated environments provide an excellent laboratory to explore this issue. HVCs are Hi clouds at radial velocities incompatible with the regular rotation of the Galactic disc15, 16, and the absorption lines of distant halo stars indicate that they are generally located several kpc above the Galactic plane (refs.17, 18, 19, 20). The overall negative velocities of HVCs (refs.16, 18, 21) suggest that they may trace the infall of early-phase low-metalicity material onto the plane of the Milky Way (refs.22, 23, 24, 25). High-Velocity Clouds (HVCs), as observed in all-sky Hi surveys such as HI4PI 21, typically appear as atomic gas streams or complexes. This is likely due to the limited resolution of these surveys ( 16′; see also Extended Data Fig. 1). A few HVCs with km/s have been observed by single-dish telescopes such as Arecibo 26 and GBT 27, as well as interferometers 28. These observations have revealed that HVCs are primarily filamentary in nature 29, 30, 31, with many filaments aligned in parallel patterns. However, interferometric observations have revealed much more complex filamentary structures within HVCs, and narrow-line cold neutral medium (CNM) components have been detected in these Hi filamentary systems 32, 33. It is possible that the strip-shaped filamentary structures observed in HVCs are influenced by the global dynamics of the gas streams or by dynamic processes within the Galactic plane 30, 34, while the transition from warm neutral medium (WNM) to CNM contributes to the formation of more complex filamentary structures 32, 33. In contrast, very-high-velocity clouds (VHVCs) are not coupled with the Galactic disc and are thought to have a different origin from typical HVCs 34, 35, 36, 37. Among them, VHVCs with km/s and located offset from the Galactic plane are clearly not co-rotating with the Galactic disk and, therefore, serve as the fiducial sample of VHVCs. Those VHVCs may represent an earlier evolutionary stage compared to HVCs and could still primarily contain WNM. Exploring the detailed inner structures of VHVCs, particularly their possible filamentary patterns, is essential for understanding the early formation of density structures in atomic clouds. However, detailed studies of VHVCs have been extremely rare in the past. Observations of the Hi 21 cm line (, =1-0; MHz) 15, with sufficiently high resolution to reveal the internal morphology and gas dynamics of VHVCs, are crucial for studying their origins and understanding the physical processes involved in the formation of hierarchical gas structures within the WNM.
@@
The HI4PI survey (refs.38, 39, 40, 41) provided so far the most detailed Hi images of all-sky HVCs21. VHVCs are dominated by the Magellanic Stream (MS) 42, 43 and the anticentre (AC) complex (refs.44, 45, 34). The VHVC component of the AC complex, denoted as G165 VHVC in this work, is the most notable one in brightness and size on the sky (Extended Data Fig. 1), which stands out as an ideal object for investigating filament formation in extreme environments. Besides the VHVC component (-330 km s-1 -250 km s-1), at the line of sight of G165 there is also an HVC component (-140 km s-1 -80 km s-1) and a foreground component (-20 km s-1 20 km s-1). G165 HVC is part of the Cohen Stream, which was measured to have a distance bracket of 5.0 to 11.7 kpc18. In contrast, the H brightness of G165 VHVC implies a distance () on the order of 15 kpc18, the value we adopted. The G165 VHVC is gravitationally unbound with a size of 100 pc, and a total gas mass of (see methods). Limited by poor angular resolution (), the inner structures of the G165 VHVC were not resolved but did show clumpy features in the HI4PI survey.
@@ A WNM system with supersonic dynamics
We mapped the G165 VHVC using the Five-hundred-meter Aperture Spherical radio Telescope (FAST), with unprecedented angular resolution (, corresponding to a spatial resolution of 13 pc at a distance of 10 kpc), velocity resolution ( km s-1), and line sensitivity (60 mK per channel). Complex inner structures of the G165 VHVC are revealed in the integrated intensity map of the Hi 21 cm line (left panel in Fig. 1). Although different velocity subcomponents of the G165 VHVC are spectrally blended, they can be well separated through careful Gaussian decomposition (Fig. 1; see Sect. Methods for details). These subcomponents are distributed in multiple layers in the position-position-velocity (ppv) cube (Fig. 2). The velocity layers are interlaced with one another and exhibit large-scale velocity fluctuations (or velocity wiggles), similar to those seen in the molecular interstellar medium 46. The central velocities () of the subcomponents can be divided into five groups ranging from km s-1 to km s-1 (Fig. 3a), though these velocity groups do not correspond one-to-one with the spatially separated gas layers. The thickness of each velocity layer is estimated to be 100 pc (Sect. Methods).
@@
The line widths (full width at half maximum; FWHM) of the velocity subcomponents are slightly larger than 20 km s-1 (Fig. 3c), the value of the typical thermal line widths of the WNM47, 48, 2. Assuming a gas kinetic temperature of the G165 VHVC of 6000 K, the typical value of WNM2, the thermal dispersion is estimated as 7 km s-1, yielding the transonic () nature of individual velocity subcomponents. However, different velocity groups are separated by 20 km s-1, suggesting the whole cloud could be in supersonic turbulence. Considering that different gas layers are intersected, forming a dynamically correlated system (see also the next sections), all velocity groups can be treated as a whole system. The total dispersion of the centroid velocities (), a good tracer of the internal turbulent motions within a cloud, is up to 20 km s-1, indicating a global of 3. We note that the observed distribution of line widths (Fig. 3c) can be well reproduced by our MHD simulation of supersonic and trans-Alfvénic gas ( and ). The same simulation also spans a large velocity range with several velocity groups (Fig. 3b), as the same in observations.
@@
The median values of column density ( cm-2) and volume density (0.2 cm-3) of the G165 VHVC (Sect. Methods) are also similar to the typical values of the WNM47. The column density probability distribution function (-PDF) of the decomposed subcomponents is higher-end suppressed (or lower-end enhanced), and deviates from a lognormal shape (Fig. 3d). The lognormal behavior of an -PDF is usually explained by supersonic turbulence49, 50, 51, but the intermittency of a supersonic system caused by head-on collisions of strongly shocked gas could lead to considerable deviations52, 53. The observed -PDF can also be depicted by our MHD simulation models of supersonic and trans-Alfvénic gas ( and ; see below).
@@
To conclude, we consider the G165 VHVC as a WNM system with supersonic dynamics. This is significantly different from the majority of the WNM in Galactic plane or in HVCs (e.g., Smith cloud), where turbulence is subsonic to transonic 14, 30, 54, 51.
@@
Filamentary structures in the G165 region, which cannot be resolved by HI4PI, are clearly seen by FAST on the Hi moment 0 maps of its VHVC, HVC, and foreground components (Extended Data Fig. 2). The HVC components show parallel filaments elongated along the Cohen Stream, similar to the famous Smith HVC 30. In contrast, both the foreground and VHVC components show web-like filamentary structures. The foreground component has no clear boundary, while the VHVC shows clear boundaries defined by long and slim filaments. Another difference between the foreground and VHVC components is that the filaments tend to be in parallel mode on the channel maps of the foreground component, but stay intersected at different velocity slices of the VHVC components (Extended Data Fig. 2). These comparisons indicate that the filamentary system of the G165 VHVC is really web-like rather than in parallel mode, which is substantially different from those in typical HVCs (e.g., Smith HVC 30) and low/intermediate velocity Hi clouds 11, 13.
@@
Filamentary structures in the G165 VHVC are visible in different layers of the velocity subcomponents (Fig. 2). However, this may be misleading due to projection effects, which enhances the surface intensity at places tangent to the line of sight. We developed a tool, ClimbRidge, based on a density ridge algorithm55, 56, to extract the density ridges of G165 VHVC in both 2D and 3D cases (Sect. Methods). ClimbRidge gives similar results as the well known tool FilFinder in 2D case (Extended Data Fig. 3), and can be naturally applied to 3D (ppv) data to trace truly velocity-coherent filaments (Fig. 2 and Extended Data Fig. 4). Minimal Spanning Trees (MSTs) of the points on the density ridges were built (Extended Data Fig. 5; see Sect. Methods), and the lengths of filament segments were measured (Fig. 3e). As one can see, the distribution of segment lengths in our MHD simulation of supersonic and trans-Alfvénic gas ( and ) remarkably resembles the observed one.
@@
The filaments in the G165 VHVC are mainly distributed in three layers in ppv space (denoted as L1, L2 and L3 in Extended Data Fig. 4). Each layer has an angular size larger than 2.5∘. On the plane of (RA, DEC), the G165 VHVC can be divided into its eastern, western, and southern parts. L1 is the main body of the western part. L2 spans from the eastern part to the western part, with a global velocity gradient of 0.2 km s-1 per arcmin (0.05 km s-1 pc-1). L3 extends from the western part to the southern part, with a global velocity gradient of 0.3 km s-1 per arcmin (0.07 km s-1 pc-1) (Extended Data Fig. 4). The flare gas above L2, with velocity spreading to km s-1, also has a filamentary structure. The filaments of the G165 VHVC are intersected with each other, building a filamentary network in the ppv space.
@@
There are long and slim filaments that go through different layers, indicating that different layers are indeed not independent and form a dynamically correlated system. Three representative filaments (denoted as B1, B2, and B3) are highlighted in the right panel of Fig. 2 and Extended Data Fig. 4. We examine these long filaments directly in the ppv cube (Extended Data Fig. 6), which reveals that their velocity structure is coherent, confirming that the density ridges trace the skeletons of the filamentary structures in ppv space. Basic parameters of the three filaments are fitted and listed in Extended Data Table 1. The velocity gradient of the filaments is comparable with that of the gas layers (0.1 km s-1 pc-1), contributing a velocity span of km s-1. This also suggests that the filaments in G165 VHVC are supersonic. The filaments of the G165 VHVC tend to be radially asymmetrical on the - plane (for example see Extended Data Fig. 6). The intensity difference between the two sides of the filament is comparable with the intensity enhancement of the central ridge (Extended Data Fig. 7). Our MHD simulations suggest that the asymmetrical density enhancement of filaments could be induced by supersonic turbulence (Fig. 3f). Although the line width analysis and simulations (Fig. 3) suggest that the majority of the components of G165 VHVC are WNM, we cannot fully exclude the presence of CNM components, as unresolved thin layers of CNM may be induced by turbulent flows57.
@@
Origins of hierarchical filamentary structures
Hierarchical filamentary structures are ubiquitous in Galactic molecular clouds (MCs), manifesting as slim curves 58, hubs 59, and webs 3. In contrast, filamentary structures in atomic clouds tend to be simpler. However, in G165 VHVC, all these filamentary types – slim curves, hubs, and webs – are present, suggesting the possibility of forming hierarchical filamentary structures without the aid of gravity, as in MCs. The web-like filaments within G165 VHVC do not merely intersect due to projection effects. Instead, different filaments intersect in the ppv space, with mass assembling at the intersection knots (see filament B1, for example, in Extended Data Fig. 6). Different velocity layers are connected by slim filaments (Extended Data Fig. 4). Along the filaments of G165 VHVC, velocity gradients comparable to those in MCs (ranging from 0.1 km s-1 pc-1 to several km s-1 pc-1) have already been established. Note that velocity gradients along filaments in MCs are common 60, 7, 8, and are often interpreted as gas flows driven by gravity. The results of this work suggest that, in addition to gravity, other mechanisms – particularly MHD turbulence – play an important role in the formation of hierarchical filamentary structures and in shaping the velocity gradients along filaments.